OpenStudy (anonymous):

Evaluate the following expression. 1/5^-2 is it 1/25?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Firstly what the question is: \[\frac{1}{5^{-2}}\] ??

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

See in this case we use the following concept: \[\large \frac{1}{x^{-a}} = x^a\] Can you do it now??

OpenStudy (anonymous):

so 1/25 is right. right??

OpenStudy (anonymous):

No.. Look the formula clearly..

OpenStudy (anonymous):

25?

OpenStudy (anonymous):

Absolutely correct.. Well done..

OpenStudy (anonymous):

thank you.. can you help me with 6^-3 too? is it 216?

OpenStudy (anonymous):

or is it 1/216?

OpenStudy (anonymous):

second one.. When you have negative powers just take the reciprocal..

OpenStudy (anonymous):

ok.. how about 153^0?? is it 1?

OpenStudy (anonymous):

Anything raised to the power is 1.. except \(0^0\) it is not defined..

OpenStudy (anonymous):

Yes it is 1...

OpenStudy (anonymous):

Anything raised to the power 0 is 1.. except \(0^0\) it is not defined..

OpenStudy (anonymous):

how about (-6)^-2? sorry...

OpenStudy (anonymous):

36?

OpenStudy (anonymous):

it can be written as: \(1/(-6)^2\)

OpenStudy (anonymous):

so is it -1/36?

OpenStudy (anonymous):

NO.. 1/36

OpenStudy (anonymous):

sorry.. real last question!! Evaluate m0 - n2 for m = 2 and n = -1. so its 2^0-(-n)^2.. i dont get this

OpenStudy (anonymous):

\(2^0 - (-1)^2 \implies 1- (-1)^2 \implies 1- (1) \implies 0\)

OpenStudy (anonymous):

huh? so its 0?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

alright thank you so much~!!!!!!!!!!!11

OpenStudy (anonymous):

Welcome dear..